Center for Astrophysical Research in Antarctica
January 24, 1999 1:00PM CST
|Chicago||AST/RO|| U of CAL-Santa Barbara |
South Pole: Bob Pernic
South Pole: Al Harper
South Pole: Bob Loewenstein
South Pole: Joe Rottman
South Pole: Jeff Peterson
- ABU is online about 3-4 weeks later than planned, but telescope is
- SPARO is on Viper and working. Because of the weather and high water vapor the
Hemp dewar will be put back on Viper for testing before the end of the season. Then
SPARO can go back on.
Transfer time of Hemp and SPARO is 10-20 minutes, which is acceptable for the
- AGO- not working as well as it should be. Waiting for 110V line as backup power
source. Bob Pernic will attempt to problem solve the continuing inefficiencies of the
- Winter helium should be fine. Supply is coming in on the supply ship. The old
dewars will be returned to McMurdo and filled there and then returned to Pole. Also
smaller dewar is coming to Pole to top off the 3000 & 1000 liter dewars once at Pole
and then fill all the other dewars that are around the base.
- AST/RO things OK. Some problems with receivers and lock boxes. Adair
sending a new lock box to Pole today. Alignment of focus still not optimized. Still have to
scramble because of helium delays.
- 3 250 liter dewars are expected by Friday for Summer helium
Project Committee Meeting To be held March 4th and 5th at
ASA wants to know planed use and needs of new 5 bay building. AMANDA will also
need to be involved in planning of Dark Sector Lab space. It was suggested that
AMANDA be invited to Yerkes for these discussions.
Budgets for year 10 are to be submitted to Nancy on or before February